OpenStudy (anonymous):

Neil started a stamp collection with 12 stamps. Every week, he adds 4 more stamps to the collection. Which function f represents the relationship between the number of stamps (s) and the number of weeks (w)? help???

OpenStudy (anonymous):

s = f(w) = 12 + 4w w = f(s) = 12s + 4 s = f(w) = 12 + 4s w = f(s) = 12w + s s = f(w) = s – 12

OpenStudy (anonymous):

someone please help i think i know but not sure

OpenStudy (anonymous):

Hi welcome to openstudy! w= # of weeks s=# of stamps you are looking for a function that represents this function (or the # of stamps) depending on the number of weeks you have. so from this you know that s=f(w). Next, you know that the constant term or the thing that can't be change is the starting number of stamps which is 12. Then, every week, you add 4 stamps, so that will be 4w Combine these things and you'll end up with........? which one do you think?

OpenStudy (anonymous):

i think the first one??

OpenStudy (anonymous):

right! :)

OpenStudy (anonymous):

awesome! thank you!

OpenStudy (anonymous):

no problem ;)

OpenStudy (anonymous):

Ryan opens a bank account with $50. Each month he deposits an amount that doubles his savings. Ryan gets $2 from the bank each month as an incentive for maintaining this savings pattern. Which recursive function represents the amount in Ryan’s account in any month?

OpenStudy (anonymous):

i got this wrong last time :(

OpenStudy (anonymous):

next = 2 × now + 2, starting at 50 next = 2 + 2, starting at 2 next = now + 2, starting at 50 next = 2 – now, starting at 50 next = 2 × now, starting at 2

OpenStudy (anonymous):

can you help please

OpenStudy (anonymous):

give me a sec. i'll try to understand it first :)

OpenStudy (anonymous):

okay thank you

OpenStudy (anonymous):

wouldnt it be the first one??

OpenStudy (anonymous):

starting at 50, doubles his savings every month, $2 from the bank each month yes, I think the first one makes more sense than the rest

OpenStudy (anonymous):

yes it makes the most sense!

OpenStudy (anonymous):

can you help on another one?

OpenStudy (anonymous):

If f(x) = 2x2 + 3x and g(x) = x – 2, what is (f + g)(2)?

OpenStudy (anonymous):

16 14 12 10 8

OpenStudy (anonymous):

f(x)+g(x) \(2x^2 + 3x + x – 2\) f(2)+g(2) \(2(2)^2 + 3(2) + 2 – 2=?\)

OpenStudy (anonymous):

14?

OpenStudy (anonymous):

yep! keep up the goodwork ;)

OpenStudy (anonymous):

i passed the quiz thank you so much!!

OpenStudy (anonymous):

i wish you could help me on this question im doing now :/

OpenStudy (anonymous):

oh is this a quiz? if so, I cannot help you because they said that we are only allowed to help other users with their assignment/projects but not with any kind of assessments (quiz, test exam) because it is considered as cheating :/ sorry. if this is assignment I can help you, if I still have more time (I have to go to my night school)

OpenStudy (anonymous):

its the tutorial to help before my big test. so it really isnt a test i should say its just helping me to pass the test once i get to it

OpenStudy (anonymous):

ok